# Variance Stabilizing Transformations for Electricity Price Forecasting in Periods of Increased Volatility

### v1

@@ 3.3 Forecast Selection and Combination Methods
Once point forecasts have been generated using multiple parameter configurations of a single variance stabilizing transformation (VST), or across several distinct VSTs, the next step is to determine which of these forecasts should be used. As demonstrated in following result Section 4, forecast performance varies significantly across markets, hours, and selected transformation parameters. Consequently, relying on a single transformation or fixed parameter may lead to suboptimal results. To address this, we propose four systematic approaches to post-process the pool of forecasts. We divide those methods into two categories: selection-based methods, which choose a single forecast from the set, and averaging-based methods, which combine multiple forecasts. In all cases. the selection of forecasts is based on mean absolute error (MAE) within a 56-day rolling or fixed window. The procedures are applied independently for each hour of a day.
@@ 3.4 Evaluation measures
where $\hat{\varepsilon}_{d,h}={p_{d,h}-\hat{p}_{d,h}}$ is the forecasting error for day $d$ and hour $h$. The aggregated measure for the whole out-of-sample period of $D$ days is defined as follows: $\text{MAE}=\sqrt{\frac{1}{D}\sum_{d=1}^{D}(\text{MAE}_{d})^{2}}$
@@ 4.1 Impact of VST parameterization
Table 1: Mean absolute error (MAE) of forecasts for each variance stabilizing transformation (VST) and each market. Columns correspond to transformations (asinh, box–cox, mlog, T-PIT) and markets (EPEX-DE, OMIE-ES, EPEX-FR). Rows report results for: (i) CB (crystal ball benchmark, using the ex-post optimal parameter for each hour), (ii) Literature (fixed parameters used in prior studies), and (iii) four proposed selection and averaging schemes ($\mathrm{SL}_{\text{fix}}$, $\mathrm{SL}_{\text{roll}}$, $\mathrm{AVG}_{\text{fix}}$, $\mathrm{AVG}_{\text{roll}}$). The best result in each column, excluding the crystal ball benchmark, is highlighted in bold.
@@ 4.2 Transformation selection and averaging
Table 2: Mean absolute errors (MAE) for individual variance-stabilizing transformations (upper subsections) and proposed selection/averaging schemes (lower subsections) across three markets (EPEX-DE, OMIE-ES, EPEX-FR) and five calibration window lengths. Results are reported separately for NARX (upper panel) and LEAR (lower panel). The best single transformation in each column is underlined, while the overall best result, including combination schemes, is shown in bold.
@@
For LEAR, T-PIT is the dominant choice, yielding the best single-transformation results in 11 out of 15 cases across markets and windows. Box-Cox and asinh are occasionally competitive, while mlog never provides the lowest errors. On the other hand T-PIT is the worst choice (underperformed even compared to model without transformation for both LEAR and NARX schemes) for the longest 1456 day calibration window. It highlights that the transformation should not only be optimized for given market but also considering the size of calibration window size.
@@ 4.3 Temporal Analysis of VST Performance
Figure 4: Improvements in forecast accuracy (absolute reduction in MAE) from variance-stabilizing transformations relative to untransformed benchmarks, reported for the rolling averaging scheme ($\mathrm{AVG}_{\text{roll}}$). Results are shown for three subperiods (2019–2020, 2021–2022, 2023–2024), five calibration window sizes, and two model classes (NARX in blue, LEAR in orange). Bars within each group correspond to calibration windows of increasing length (56, 182, 364, 728, 1456 days).
@@ 5 Conclusions
This study examined the role of variancestabilizing transformations (VSTs) in day-ahead electricity price forecasting across three liberalized markets—Germany, Spain, and France—over the period 2015–2024. By analyzing four widely used transformations (asinh, Box–Cox, mirror logarithm, and T-PIT) within two distinct model classes (NARX and LEAR), we show that the choice of transformation and its parameterization critically influences forecast accuracy, with no single transformation performs best across all markets, horizons, and calibration windows

### v2

@@ 3.3 Forecast Selection and Combination Methods
Once point forecasts have been generated using multiple parameter configurations of a single variance stabilizing transformation (VST), or across several distinct VSTs, the next step is to determine which of these forecasts should be used. Relying on a single transformation or fixed parameter may lead to suboptimal results. To address this, we propose four systematic approaches to post-process the pool of forecasts. We divide those methods into two categories: selection-based methods, which choose a single forecast from the set, and averaging-based methods, which combine multiple forecasts. In all cases. the selection of forecasts is based on mean absolute error (MAE) within a 56-day rolling or fixed window. The procedures are applied independently for each hour of a day.
It is important to note that the proposed procedures do not involve direct optimization or tuning of VST parameters within the forecasting models. Instead, multiple parameter configurations are considered in parallel, and the final forecast is selected or combined ex post based on recent forecasting performance, which constitutes the main contribution of this study.
@@ 3.4 Evaluation measures
where $\hat{\varepsilon}_{d,h}={p_{d,h}-\hat{p}_{d,h}}$ is the forecasting error for day $d$ and hour $h$. The aggregated measure for the whole out-of-sample period of $D$ days is defined as follows: $\text{MAE}=\sqrt{\frac{1}{D}\sum_{d=1}^{D}(\text{MAE}_{d})^{2}}$. All MAE values are expressed in EUR/MWh.
@@
The aim of performing the CPA test is to statistically assess whether differences in forecast accuracy between two competing models are significant over time. In particular, it tests the null hypothesis of equal predictive performance by accounting for temporal dependence and heteroskedasticity in the loss differential series. A small p-value indicates that the null hypothesis can be rejected, implying that one model exhibits statistically different (superior or inferior) predictive performance relative to the other. The CPA test is applied pairwise to compare selected forecasting configurations and is used to support relative model assessment rather than as a simultaneous multiple-hypothesis testing procedure.
In this paper, we focus on accuracy measures aggregated for all 24 hours. However an hour-by-hour accuracy analysis was also performed. Due to space limitations and the large number of models and configurations considered, these results are not included in the manuscript but are available from the author upon request.
@@ 4.1 Impact of VST parameterization
Table 1: Mean absolute error (MAE) of forecasts for each variance stabilizing transformation (VST) and each market. Rows report results for: (i) CB (crystal ball benchmark, using the ex-post optimal parameter for each hour), (ii) Literature (fixed parameters used in prior studies), and (iii) four proposed selection and averaging schemes. The best result in each column, excluding the crystal ball benchmark, is highlighted in bold.
@@ 4.2 Transformation selection and averaging
Table 2: Mean absolute errors (MAE) for individual variance-stabilizing transformations (upper subsections) and proposed selection/averaging schemes (lower subsections) across five calibration window lengths. The best single transformation in each column is underlined, while the overall best result, including combination schemes, is shown in bold.
@@
For LEAR, T-PIT is the dominant choice, yielding the best single transformation results in 11 out of 15 cases across markets and windows. Box-Cox and asinh are occasionally competitive, while mlog never provides the lowest errors. On the other hand T-PIT is the worst choice (underperformed even compared to model without transformation for both LEAR and NARX schemes) for the longest 1456 day calibration window. It highlights that the transformation should not only be optimized for given market but also considering the size of calibration window size.
@@
Figure 4: Average MAE improvements relative to the baseline model without variance-stabilizing transformation for three electricity markets and two base models (NARX and LEAR).
To facilitate a direct comparison of the effects of adaptive parameterization and transformation selection, Figure 4 reports relative improvements in forecasting accuracy with respect to the baseline model without variance-stabilizing transformation. For each configuration, the improvement is computed as $\Delta_{\mathrm{MAE}}=\frac{\text{MAE}_{\mathrm{id}}-\text{MAE}_{i}}{\text{MAE}_{\mathrm{id}}}\times 100\%,$ where $\text{MAE}_{\mathrm{id}}$ denotes the MAE of the model estimated on untransformed prices and $\text{MAE}_{i}$ denotes the MAE obtained for a given variance-stabilizing transformation and selection scheme.
To be more specific the baseline level represents the average MAE of models estimated without variance-stabilizing transformation, averaged over all calibration window sizes. The bars labeled ’Literature’ and ’Parameter optimization’ report average MAE values obtained using variance-stabilizing transformations, averaged jointly over all considered transformations and calibration window sizes, with parameters selected either using fixed literature-based rules or adaptively using the rolling $\mathrm{AVG}_{\text{roll}}$ scheme, respectively. Finally, the bar labeled ’Transformation optimization’ reports average MAE values averaged over all calibration window sizes for models in which the variance-stabilizing transformation itself is selected adaptively using the $\mathrm{AVG}_{\text{roll}}$ criterion.
The results confirm that static, literature-based parameter choices only yield limited improvements over the baseline, and in some cases are slightly worse. In contrast, adaptive parameter optimization leads to systematic gains across markets and base models, which increase further when the transformation itself is also selected adaptively. Overall, adaptive transformation selection provides the greatest and most consistent improvements in accuracy. These effects are more pronounced for the LEAR model than for NARX, indicating that the benefits of variance-stabilizing transformations are expected to weaken for more complex models and may diminish for highly flexible nonlinear structures such as deep neural networks (lag:mar:des:wer:21) or hybrid and attention-based approaches (ma:mei:22).
@@ 4.3 Temporal Analysis of VST Performance
Figure 5: Improvements in forecast accuracy from variance stabilizing transformations relative to untransformed benchmarks, reported for the rolling averaging scheme ($\mathrm{AVG}_{\text{roll}}$). Results are shown for three subperiods (2019–2020, 2021–2022, 2023–2024), five calibration window sizes, and two model classes (NARX in blue, LEAR in orange).
@@ 5 Conclusions
This study examined the role of variance stabilizing transformations (VSTs) in day-ahead electricity price forecasting across three liberalized markets – Germany, Spain, and France –over the period 2015–2024. By analyzing four widely used transformations (asinh, Box–Cox, mirror logarithm, and T-PIT) within two distinct model classes (NARX and LEAR), we show that the choice of transformation and its parameterization critically influences forecast accuracy, with no single transformation performs best across all markets, horizons, and calibration windows
@@ Acknowledgements
Declaration of generative AI
During the preparation of this work the author used ChatGPT (Open AI) in order to language editing. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.